I cannot understand why the formula of net present value is the following

rn に0

why

and not

?
Answer #1

let me know if you need any clarification ...

Explanation: Present value is the value of money today which will be received in future.
Money has the capacity to earn interest and therefore same money today worth more than received in future.
Therefore each year we have to divide the cash flow with the interest rate^ times (in year) to convert this in present value term.
Example:
If we have $ 100 receivable for 5 year at the end of each year we need to convert this into present value. We have to use below formula
Assuming interest rate is 10%
Year Cash flow present value
1 100 =100/(1.1)^1 90.90909
2 100 =100/(1.1)^2 82.64463
3 100 =100/(1.1)^3 75.13148
4 100 =100/(1.1)^4 68.30135
5 100 =100/(1.1)^5 62.09213